⟹ ⟹ a^{2} + b^{2} = c^{2}

\begin{equation}
f(x)=\sum_{n=0}^{\infty}\frac{f^{(n)}(x)}{n!}
\end{equation}

f(x) = ∑_{n=0}^{∞} \frac{f^{(n)}(x)}{n!}

latex  foo.tex produces foo.dvi (dvi file)
pdflatex  foo.tex produces foo.pdf (pdf file)
dvips  -o  foo.ps  foo.dvi produces foo.ps (postscript file)

latex  foo.tex
dvips  -o  foo.ps  foo.dvi
ps2pdf  foo.ps produces foo.pdf (pdf file)

htlatex  foo.tex

mk4ht  oolatex  foo.tex (Linux/Mac platforms)
oolatex  foo.tex  (Windows platform)

mk4ht  mzlatex  foo.tex (Linux/Mac platforms)
mzlatex  foo.tex  (Windows platform)

mk4ht  mzlatex  foo.tex  "html,mathplayer" (Linux/Mac platforms)
mzlatex  foo.tex  "html,mathplayer"  (Windows platform)

mk4ht  mzlatex  foo.tex  "html,pmathml" (Linux/Mac platforms)
mzlatex  foo.tex  "html,pmathml"  (Windows platform)

mk4ht  mzlatex  foo.tex  "html,mathplayer,imgdir:images/" (Linux/Mac platforms)
mzlatex  foo.tex  "html,mathplayer,imgdir:images/"  (Windows platform)

mk4ht  mzlatex  foo.tex  "html,pmathml,imgdir:images/" (Linux/Mac platforms)
mzlatex  foo.tex  "html,pmathml,imgdir:images/"  (Windows platform)

mk4ht  mzlatex  foo.tex  "html,pmathml" (Linux/Mac platforms)
mzlatex  foo.tex  "html,pmathml"  (Windows platform)